PVDIS Run Plan Oct-Dec, 2009
DRAFT
last updated: 11/06/09, 2:38pm
General Information: draft calendar.
Go to today's run plan.
Things to do during commissioning (in preliminary order), including person in charge, [estimated beam time], and planned time:
- Day 1, 11/3, Tuesday, day shift: continue detector installation
- Day 1, 11/3, Tuesday, swing shift:
- All shift workers on call at the beginning of shift.
- Continue installation until near 8pm. RC inform the MCC we will be ready for tuned beam by 8pm.
- Make sure both detector stacks are push in the hut and the hut doors are closed.
- Close up the Hall by 8pm, prepare for beam tune. Call in the shift leader for the swing shift. Target MUST BE in empty position throughout the shift.
- beam startup using EMPTY target. Check the following:
- Beam energy lock: ON Hall A.
- Beam entrance and exit window cover OUT.
- Beam tune straight, not through Compton magnets.
- Verify settings of target Ion chamber trip limit (MCC with Radcon).
- Record BPM A and B readings in (x, y).
- ?? Ask for harp scan, require that beam $\sigma_x$: 100 - 300microns, $\sigma_y$:100 - 300microns on two hall A harps.
- Compton initial tune - Alexandre with MCC optics experts;
- As soon as we have stable CW beam, do PITA scan (part of helicity board checking) - Rupesh.
- Day 2, 11/4, Wednesday owl shift:
- Shift leader is required. Other shift workers on call.
- Continue with Compton tune and PITA scan (part of helicity board checking) - Alexandre, Rupesh.
- At 6am, if there is good CW beam in the hall, page Sasha Glamazdin for initial Moller commissioning [1-2 hours]
- Prepare for access after Moller is done, near 8am (restricted or controlled???)
- Day 2, 11/4, Wednesday, day shift
- Hall should be open throughout the shift.
- Target cool down - Dave Meekins. Target operator is required from now on. TO check with Jian-ping to get a refresh training.
- Techs continue installing the detectors on the Right arm;
- PVDIS crew work on cabling up the Right arm, do quick cosmic checks.
- Day 2, 11/4, Wednesday, swing, and Day 3, 11/5, Thursday, owl shifts.
- PVDIS crew continue installing cables and detector checks.
- Finish replacing S1,S2 PMTs - Bogdan.
- Check with target expert and Kees to make sure it's okay to put beam on target.
- After installation is finished, close up the hall and prepare for beam.
- Set HRS magnets to the desired momentum, and angles.
- Survey spectrometer angles.
- If this is before 10pm, do the following
- Compton tune - Alexandre
- Compton electron detector work - Michel.
- If this is after 10pm, call in shift leader and do the following (TO should be on shift by default, 3rd person remains on call):
- Start full beam recovery procedure. [1-2 hours] - Jin (SL), can call in an experienced person for help.
- detector
check at low current using VDCs and HRS runs. Make sure all ADC and TDC
signals are okay (S1, S2, GC, Shower, Preshower, VDC), on optics target - Vince, Jin. [2 hours]
- establish
PS, SH, GC, S1, S2 high voltage (gain-matching to make sure raw ADC
peaks are at about the same place), on optics target - Vince, Jin. [maybe another 1-2 hours]
- This list should take the whole owl shift. But, if there is more time, here is what else we can do:
- Q2 measurement at low current (<=1uA) on LD2 target. - Dustin
- The only run on cryotarget for this owl shift is for the Q2 measurement. Use solid targets for other studies!
- TO for owl shift? Rupesh.
- Day 3, 11/5, Thursday, day shift
- Shift leader and TO are required. 3rd person on call.
- MCC
is planning on an injector work for about 4 hours. This is try to get
to 100uA.
- Meantime, techs work on RHRS Q2. PVDIS crew work on triggers and lead glass cables.
- Close up hall by 2:30pm, first attempt on Compton, commission photon and electron detectors - Diancheng, Alexandre, Megan. [up to 4 hours]
- PVDIS crew work on DAQ checks, analyze data from owl shifts and figure out cable length adjustment.
- Day 3, 11/5, Thursday, swing shift. Shift crew: John LeRose, Jin Huang, John Arrington
- commission
cryotarget: attemp a high current on LD2, up to 120uA, find out if the target
(cooling and high power heater) can take it. Production run current will be 80~90% of maximum - Jian-Ping. [1-2 hours]
- establish luminosity limits (target, radiation, source), can be done at the same time as previous step:
- establish Lumi high voltage for currents up to 100uA (or highest possible). This should be checked again when - Luis?, Dustin. [2 hours]
- Right arm detector checkout - Vince
- detector
check at low current using VDCs and HRS runs. Make sure all ADC and TDC
signals are okay (S1, S2, GC, Shower, Preshower, VDC), on optics target - Vince, Jin. [2 hours]
- establish
PS, SH, GC, S1, S2 high voltage (gain-matching to make sure raw ADC
peaks are at about the same place), on optics target - Vince, Jin. [maybe another 1-2 hours]
- Adjust HV to make sure no ADC "rolls over".
- Day 4, 11/6, Friday, owl shift. Shift crew: Zafar Ahmed, Robert Michaels, Li Ye
- low current Q2 measurement - Bob.
- boiling test (short version, complete as much as possible), establish running condition for LD2 - Diancheng, Xiaoyan, Luis or Rupesh will stay around for the first a few runs. Width results should be posted in halog. by the end of shift [total 4-8 hours, or until end of shift].
- Day 4, 11/6, Friday, day shift. Shift crew: Elena Long, Lamiaa El Fassi, Nuruzzaman
- Injector work to increase max current into hall [1 hour]
- Controlled access to fix Right Q2, could take up to 8 hours.
- Also add more backup ADCs - Bob.
- Fix left arm lead glass cabling
- Fix numerous stuff on right arm
- Day 4, 11/6, Friday, swing shift. Shift crew: Jin Huang, Bo Zhao, John Arrington
- Moller measurement. This should also produce a sign check for the new helicity
board - (Sasha Glamazdin) [4 hours, day time preferred]
- If all magnets are up:
- Q2 measurement at low current, both arms - Jin, John [1 hour, low current]
- Possible accessed to fix PVDIS DAQ [2 hours]
- Finish boiling test, (short version), establish running condition for LD2 - Diancheng, Xiaoyan. [total 8/2=4 hours]
- For
the weekend, if it is expected that there will be a couple of hours and
we cannot do production or other commissioning work, page Bob Michaels
and he will come up with a plan.
- Day 5, 11/7, Saturday, owl shift. Shift crew: David Flay, TO?, Rupesh Silwal
- Dithering magnet checkout - Rupesh?, Dustin [1-2 hour]
- continue with boiling test, (short version), establish running condition for LD2 - Diancheng, Xiaoyan.
- Day 5, 11/7, Saturday, day shift. Shift crew: Alexandre Deur, Lamiaa El Fassi, Brian Quinn
- Access for PVDIS DAQ tuning.
- Compton tuning starting at 3pm.
- Day 5, 11/7, Saturday, swing shift. Shift crew: Ole Hansen, Jin Huang, John Arrington
- Compton tuning until 6pm. In parallel, Right arm detector commissioning;
- Compton electron detector work.
- Possible beam position scan on LD2 cell.
- Continue tuning PVDIS DAQ, multiple accesses possible.
- Day 6, 11/8, Sunday, owl shift. Shift crew: Zafar Ahmed, TO?, Rupesh Silwal.
- possible production, or continue tuning PVDIS DAQ.
- Day 6, 11/8, Sunday, day shift. Shift crew: Brian Quinn, Lamiaa El Fassi, Li Ye
- Day 6, 11/8, Sunday, swing shift. Shift crew: Bo Zhao, Jin Huang, Serkan Golge.
- Production on right arm with accesses to fix PVDIS DAQ.
- Day 7, 11/9, Monday, owl shift. Shift crew: SL?, TO?, Rupesh Silwal
- Shift started with access to fix PVDIS DAQ, then production on right arm only.
- Day 7, 11/9, Monday, day shift. Shift crew: Jian-ping Chen, Robert Michaels, Nuruzzaman
- Day shift spent on controlled access to fix left dipole and various fix/tunes on PVDIS DAQ.
- Day 7, 11/9, Monday, swing shift. Shift crew: Alexandre Deur, Jin Huang, Serkan Golge.
- Compton tune, for as long as they need, or until 1am
- Bull's eye scan on LD2 cell.
- Production.
- Day 8, 11/10, Tuesday, owl shift.
- Day 8, 11/10, Tuesday, day shift.
- Compton tune until 10am
- Access to fix left dipole.
- Production.
- Day 8, 11/10, Tuesday, swing shift.
- by this time, beam charge feedback should be checked and working, HCBA qualification - Kent?.
- Meantime, accomodate Compton edet work - Michel, Charles. [1 hour per day]
- Tasks with "*" means it requires relatively long time and should be seperated to several times.
- The following are copied from H-III runplan and should be studied
- IPM cavity triplet high-current commissioning [2] [3]
- IPM cavity triplet nanoamp commissioning [2] [3]
- Make sure target chamber is good enough for beam tune (RC check with Ed).
To do: MCC nulls the vertical polarization [2 hours]To do: Mini-spin dance to tune wien angle. Require coordinate with MCC and Hall B in advance - (Sasha) [3 points, up to 6 hours]To do: Moller commissioning (injector work) - Sasha with injector people [3 hours]To do: fix carbon optics position???To do: bcm calibration (invasive to Hall B, not coordinated yet!) for both HAPPEX copies and PVDIS copies (brief version?), with empty target - Kai, Rupesh, Xiaoyan [time? 2-4 hours? Must be done when the beam is relatively stable]If the DAQ is working:- bcm linearity/deadtime study (off-line?) - can be combined with bcm calibration.
- *tagger test: current scan with LD2 target, do online analysis of tagger counts - Kai, Xiaochao. [need taggers to work, 1-2 hour]
Things to do periodically:- finish full target boiling test within the first week - Diancheng, Xiaoyan; [could split into several separate tests. time?]
- beam half-waveplate change every two days.
- Q2 measurement: run HRS runs with VDCs on at low currents -- >= 1 time a week
- low current measurement on both spectrometers, with VDCs ON: twice a day (for PID).
- Dummy Al target asymmetry measurement (follow time posted on calendar).
- Moller
measurement -- every week and with each major change to the
accelerator. Prefer two back-to-back Moller measurements before and
after beam half-waveplate change. Could be more often if Compton does
not work well.
- Beam energy measurement (ARC) -- every week and with each major change to the accelerator (or once only for each pass?).
- bcm
linearity/deadtime study: run at different currents with 10uA steps,
other conditions same as production. -- 2-3 times a week (or more
often?)
- bcm calibration: every two weeks (or less often?)
- BPM fixed-gain calibration (not less than 1 in
three weeks) [follow procedure for Bull's eye scan].
- Bull's eye scan (2 times) [procedure given below]. ???
- Need to do spectrometer detector test by
turning ON detectors including VDCs for calibration purpose (be
specific which calibration?). [make a detail plan].
Things to do once during the whole period:
- mini-spin dance to check transverse polarization (as needed)???
- Rescattering radiative tail cross-section study ???
- Transverse asymmetry measurement (invasive to Hall B). It will take one shift to setup the vertical polarization. Hall B also want a pass change at the same time (another shift).
Expert list:
- Feedback, injector setup, dithering, beam systematics: Kent Paschke (lead), Mark Dalton, Rupesh Silwal
- HRS Detectors and PID analysis: Vince Sulkosky (lead), Jin Huang, plus our students & Ramesh
- HRS Optics: Nilanga, Kiad
- HRS kinematics (Qsq), backgrounds: Dustin McNulty, plus our students & Ramesh.
- HAPPEX DAQ: Luis Mercado, Rupesh Silwal, plus our students
Overall:
- Run plans are executed by the Run Coordinator.
- All changes need explicit approval from the Run Coordinator.
- Run coordinator must clear with the spokespeople for any major change.
What we should keep an eye on:
- Make sure spectrometers and target are surveyed before the commissioning starts.
- Make sure all online replay are in place and all instructions for shift workers printed out and placed at obvious places;
- Check start-of-run and end-of-run CODA log entry, make sure the comment and other info are ok.
- Remember to cycle quads (Q2 and Q3 only) when going from
lower momentum to higher momentum (no need to cycle them when going
from higher to lower momentum).
Meetings:
- Organize daily counting house 2nd floor meeting (3 PM) to discuss:
- major problems during the last three shifts;
- completed part of the run plan;
- run plan of the next day. Next major changes.
- Analysis meeting (daily) at counting house 2nd floor (3:30pm):
- status of online software, new variables to add, major changes.
- ongoing calibration, new tasks of data analysis.
- data quality control, update of online standard histograms.
- WAC reports, include WAC weblink here.
- expert-shifts for data quality control.